# On misconceptions about the Brier score in binary prediction models

### v3

@@ Abstract.
The Brier score is a widely used metric evaluating overall performance of predictions for binary outcome probabilities in clinical research. However, its interpretation can be complex, as it does not align with commonly taught concepts in medical statistics. Consequently, the Brier score is often misinterpreted, sometimes to a significant extent, a fact that has not been adequately addressed in the literature. This commentary aims to explore prevalent misconceptions surrounding the Brier score and elucidate the reasons these interpretations are incorrect.
@@ 1. Introduction: What is the Brier score?
Given a set of probabilistic predictions $p_{i}$ and corresponding observed outcomes $y_{i}$, the Brier score is defined as:
@@
Typically, the $y_{i}$ in such prediction models are assumed to be realisations of independent Bernoulli random variables $Y_{i}\sim Bern(q_{i})$, where $q_{i}\in[0,1]$ [22, 33]. Simplifying, one can think of a Bernoulli variable like a potentially biased coin flip: each patient either has the event 1 or 0, based on their individual risk. Two random variables are independent if the probability of an outcome for one is unaffected by the outcome of the other, for a formal explanation see, e.g., [41]. Correspondingly the best i-th prediction that can be obtained is the true underlying probability, i.e, $p_{i}=q_{i}$.
Brier score offers a comprehensive evaluation of probabilistic predictions and is a strictly proper scoring rule [9, Theorem 1], meaning that it incentivizes honest probabilistic forecasting as in expectation it is minimised if and only if the predictions are the true probabilities [20]. Note that there is a difference between good predictions with respect to true probabilities versus usefulness of the predictions in a clinical context: While a clinician might find it easiest to work with a classification of patients into zeros and ones or something very close to that, the quality of a probabilistic prediction model is judged based on how close the predictions are to the true probabilities. As the Brier score is strictly proper, the expectation of Brier score is minimised if and only if the predictions correspond to the true probabilities. If desired, based on the predictions, as a second step the predictions can be used to derive and evaluate a classification [21] or analysed with respect clinical consequences, e.g., via net benefit [36]. However, the qualities such evaluations judge is different to the Brier score [5, 21], in particular these are not strictly proper.
@@
Despite its widespread use [37, 28, 31], the Brier score is potentially often misunderstood in clinical research. Unlike more familiar statistical notions, it does not fit neatly into traditional statistical concepts potentially commonly taught in medical education [25, 4, 30, 32]. Furthermore, the Brier score is mathematically equivalent to the mean squared error (MSE), a concept introduced by C.F. Gauss [18]. The MSE is widely applied in ordinary least squares regression [12, §$\S$ 11.3.1], statistical learning [26, §$\S$ 2.2.1], or machine learning evaluation [15]. The connection of Brier score and MSE can also lead to confusion. Brier score and MSE are used in different contexts, as the Brier score compares a probability to an outcome of the binary random variable in the sense of scoring rules [20], while the MSE usually compares two real continuous values, in statistics typically comparing an estimator to the true value [12]. In particular, misconceptions about the Brier score are not uncommon and can sometimes be reinforced by potentially misleading statements in the literature [35, 36, 34, 1, 40, 11]. Given the importance of accurate interpretation in clinical applications, it is crucial to address these misunderstandings.
Several works have explored the evaluation of binary predictive models. [36] thoroughly reviews recent and more traditional measures for prediction models, Harrell [22, §$\S$ 10] provides a comprehensive overview of regression model performance metrics, Steyerberg [34] offers far-reaching guide to prediction model building in medicine, while [42] discusses the Gini coefficient, Pietra index, and other classification measures. Other studies, such as [24], examine alternative metrics designed to improve model performance assessment, and [6] compares the Brier score with net benefit analysis in diagnostic testing. Furthermore there are various decompositions of the Brier score [43], some can also be used to construct statistical tests for the Brier score [33]. Brier score also remains significant in AI-driven medical predictions [10], and can also be used for evaluation of survival outcomes [1].
@@
The evaluation of binary prediction models has been extensively addressed in the medical literature. Steyerberg et al.[36] provide a comprehensive review of traditional and more recent performance measures for clinical prediction models. Harrell offers a detailed discussion of regression model evaluation metrics [22]. Steyerberg also presents a thorough guide to the development and validation of clinical prediction models [34]. Wu and Lee discuss alternative metrics such as the Gini coefficient and Pietra index [42]. Other studies have focused on expanding the toolbox for performance assessment, including alternative metrics tailored to clinical utility [24]. Assel et al. compared the Brier score with decision-analytic approaches like net benefit analysis in diagnostic test evaluation [6].
Further, various decompositions of the Brier score have been proposed to better understand its components [43, 33]. The Brier score continues to be an important metric in AI-based predictive modeling in healthcare[10], and applies equally to the evaluation of prognostic models for survival outcomes [1].
@@
This commentary aims to clarify common misunderstandings about the Brier score, explain why they arise, and provide guidance on its appropriate interpretation in clinical research. As it is also directed at researchers in medicine, we first introduce some notions, after which we give a summary of the misconceptions.
@@
  • Perfect prediction: In clinical prediction models, a perfect prediction means that the predicted probability of an outcome exactly matches the true underlying risk for each individual patient. In the case of our setting of the Brier score, if the true probabilities are given by $q=(q_{1},\cdots,q_{n})$, the perfect prediction is given by $p=q$.
    Suppose a model estimates that a patient has a 20% chance of being readmitted to the hospital within 30 days. If that patient’s true risk is also exactly 20%, then the model has made a perfect prediction for that individual.
Quick reference for practitioners for the misconceptions
@@
# 1: Brier score of 0 = perfect model A Brier score of 0 implies extreme predictions (0% or 100%) that exactly match outcomes. This is odd in practice and may indicate overfitting or an error.
@@
# 4: A Brier score near $\bar{y}-\bar{y}^{2}$ means the model is useless Even perfect predictions can yield a Brier score near $\bar{y}-\bar{y}^{2}$ if the true risks are close to the mean incidence. This does not imply non-informativeness.
# 5: Brier score cannot exceed $\bar{y}-\bar{y}^{2}$ for reasonable predictions This is a common belief, but in small samples and some underlying distribution, the Brier score can exceed the threshold due to chance or reasonable but not perfect predictions.
@@
Some of our points have been previously observed. We review related references that observe similar findings. However, given the widespread use of the Brier score, our literature review is necessarily partial. [27] outlines examples where the model comparison of expected BSs is potentially contrary to how a human would judge. [33] outlined the distinction between calibration and prediction errorand the misunderstanding that lower Brier score indicates better calibration.
@@ Acknowledgements
2. More on the BS
@@
In this section, we review key properties of the BS, delve into some mathematical points, and discuss the means we will use to understand the BS.
As a first point, when considering the Brier score of (1.1), there are two simple related BSs. On the one hand, one can compare (1.1) to the Brier score when entering $p_{1/2}=(1/2,\cdots,1/2)$, giving $BS(p_{1/2},y)=1/4$. On the other hand, we can compare (1.1) to the Brier score when entering the mean incidence $\bar{y}_{v}=(\bar{y},\cdots,\bar{y})$, which gives
@@
Clearly, $BS(p_{1/2},y)=1/4\geq BS(\bar{y}_{v},y)=\bar{y}-\bar{y}^{2}$. Both for very high or very low incidences, $BS(\bar{y}_{v},y)$ is low. Furthermore we note that $\bar{y}-\bar{y}^{2}$ si symmetric around $0.5$, see Figure 1, and e.g. for $\bar{y}=0.1$ or $\bar{y}=0.9$, $BS(\bar{y}_{v},y)=0.09$.
2.1. Key properties about the BS
We summarise essential insights into the Brier score.

### v4

@@ Abstract.
The Brier score is a widely used metric evaluating overall performance of probabilistic predictions for binary outcomes in clinical research. However, its interpretation can be complex, as it does not align with commonly taught concepts in medical statistics. Consequently, the Brier score is often misinterpreted, sometimes to a significant extent, a fact that has not been adequately addressed in the literature. We aim to explore prevalent misconceptions surrounding the Brier score and elucidate the understanding and interpretation of Brier scores for practitioners.
@@ 1. Introduction: What is the Brier score?
Given probabilistic predictions $p_{i}$ and observed outcomes $y_{i}$, the Brier score is defined as:
@@
Typically, the $y_{i}$ in such prediction models are assumed to be realisations of independent Bernoulli random variables $Y_{i}\sim Bern(q_{i})$, where $q_{i}\in[0,1]$ [21, 34]. A Bernoulli random variable is like a potentially biased coin flip: each patient either has the event 1 or 0, based on their individual risk. Two random variables are independent if the probability of an outcome for one is unaffected by the outcome of the other [42]. Correspondingly the best i-th prediction that can be obtained is the true underlying probability, i.e, $p_{i}=q_{i}$.
Brier scores offer a comprehensive evaluation of probabilistic predictions and are strictly proper [8, Theorem 1], meaning that in expectation it is minimised if and only if the predictions are the true probabilities [19]. Note the distinction between accurate probabilistic predictions and clinical usefulness: while clinicians may prefer binary classifications, the quality of a probability prediction is judged by how closely its probabilities reflect the true risks. The Brier score is an evaluation measure that quantifies this closeness of predictions to true risks. If desired, predicted probabilities can be assessed for clinical impact, e.g., via net benefit [37], or used to derive a classification [20].
@@
Despite its widespread use [38, 28, 32], the Brier score is potentially often misunderstood in clinical research. Unlike more familiar statistical notions, it does not fit neatly into traditional statistical concepts potentially commonly taught in medical education [25, 4, 30, 33]. Furthermore, the Brier score is mathematically equivalent to the mean squared error (MSE), a concept introduced by C.F. Gauss [17]. The MSE is widely applied in ordinary least squares regression [11, §$\S$ 11.3.1], statistical learning [26, §$\S$ 2.2.1], or machine learning evaluation [14]. The connection of Brier score and MSE can also lead to confusion. Brier score and MSE are used in different contexts, as the Brier score compares a probability to an outcome of the binary random variable in the sense of scoring rules [19], while the MSE usually compares two real continuous values, in statistics typically comparing an estimator to the true value [11]. In particular, misconceptions about the Brier score are not uncommon and can sometimes be reinforced by potentially misleading statements in the literature [36, 37, 35, 1, 41, 10]. Given the importance of accurate interpretation in clinical applications, it is crucial to address these misunderstandings.
@@
Evaluation of binary prediction models has been widely studied in the medical literature. A review of traditional and modern performance measures is given in [37], which are also discussed in Harrells book [21, §$\S$ 10] for regression models. Steyerberg’s book [35] offers a comprehensive guide to clinical model development and validation. Alternative classification metrics such as the Gini coefficient and Pietra index were examined in [43]. Other works explore alternative evaluation scores [23] and compare metrics like the Brier score with net benefit analysis [5]. The Brier score has also been decomposed for deeper insights [44], and remains relevant in AI-based medical prediction [9] or survival outcome evaluation [1].
@@
We aim to clarify common misunderstandings about the Brier score, explain why they arise, and provide guidance on its appropriate interpretation in clinical research. As it is also directed at researchers in medicine, we first introduce some notions, and give a summary of the misconceptions. Then we go through the misconceptions and end with a conclusion. Supplementary analysis is kept in the Appendix for the interested reader.
@@
  • Perfect prediction: In clinical prediction models, a perfect prediction means that the predicted probabilities of outcomes exactly match the true underlying risk for each individual patient. In the case of our setting of the Brier score, if the true probabilities are given by $q=(q_{1},\cdots,q_{n})$, the perfect prediction is given by $p=(p_{1},\cdots,p_{n})=(q_{1},\cdots,q_{n})$.
    As an example, suppose a model estimates that a patient has a 20% chance of being readmitted to the hospital within 30 days. If that patient’s true risk is exactly 20%, then the model has made a perfect prediction for that individual.
Quick reference on common misconceptions
@@
# 1: Brier score of 0 = perfect model A Brier score of 0 implies extreme predictions (0% or 100%) that exactly match outcomes. This is odd in practice and typically indicates errors.
@@
# 4: A Brier score near $\bar{y}-\bar{y}^{2}$ means the model is useless Even perfect predictions can yield a Brier score near $\bar{y}-\bar{y}^{2}$ if the true risks are close to the mean incidence. This does not necessarily imply non-informativeness.
# 5: Brier score cannot exceed $\bar{y}-\bar{y}^{2}$ for reasonable predictions As a realisation of a random variable, the Brier score can exceed the threshold due to chance or reasonable predictions.
@@
Some of our points have been previously observed. We review related references that observe similar findings. However, given the widespread use of the Brier score, our literature review is necessarily partial. [27] outlines examples where the model comparison of expected BSs is potentially contrary to how a human would judge. [34] outlined the distinction between calibration and prediction error, emphasizing the misunderstanding that low Brier score indicates good calibration, while [24] notes that Brier score comparisons across datasets should be avoided as it depends on the incidence rate.
@@ Acknowledgements
2. Main properties of the Brier score
The Brier score in (1.1) is a measure to quantify the accuracy of probabilistic predictions, taking values between 0 and 1 with lower values indicating better performance. As the $y_{i}$ are realisations of random variables, the Brier score is a random variable. Hence any evaluation of the Brier score has a random component. It is strictly proper, meaning in expectations the perfect prediction uniquely minimises the Brier score of (1.1). Even more holds: in expectation the Brier score preserves the Euclidean distance order $l_{2}$ between predictions $p\in[0,1]^{n}$ to the true probability $q\in[0,1]^{n}$, where $l_{2}(p,q)=\sqrt{\sum_{i=1}^{n}(p_{i}-q_{i})^{2}}$ [31].
Suppose we have two sets of predicted probabilities, $p$ and $p^{\prime}$, from two different models, and let $q$ represent the true probabilities. If $p$ is closer to $q$ than $p^{\prime}$ is in terms of Euclidean distance, i.e.,
$l_{2}(p,q)<l_{2}(p^{\prime},q),$
then the expected Brier score for $p$ will be lower than that of $p^{\prime}$.
To build intuition, consider a 2-dimensional example where the true probabilities are $q=(1/2,1/2)$. In this case, any prediction falling inside a circle centered at $(1/2,1/2)$ will have a lower expected Brier score than one on or outside the circle. This provides a geometric view of probabilistic accuracy: the closer your predictions are to the truth (in Euclidean distance), the better the model performs.
$x$$y$111111110.20.20.20.2(12,$(\frac{1}{2},\frac{1}{2})$
We can further quantify the expected behavior when slightly perturbing the predicted probabilities away from the true values by $\varepsilon$. This results in an expected Brier score increase of $\varepsilon^{2}$. To be more specific, if, say, the predicted probabilities are $0.1$ more off, the expected Brier score will be $0.010.1^{2}=$ 0.01 bigger. In contrast, changes in the true probability can lead to more significant shifts in expectation.
While individual true probabilities are unobservable in practice, the observed prevalence provides an estimate of the expected prevalence and can serve as a non-informative reference point for comparison. For idealised scenarios, the law of large numbers or central limit theorem can be used to argue that the observed score is a reliable estimate of its expected value. For large datasets, this can be used to justify treating the empirical Brier score as a stable summary measure of model performance. To summarise and simplify our previous points, we note that an observed Brier score is a function of
  1. (I)
    the underlying true probabilities (the $q_{i}$s),
  2. (II)
  3. (III)
The influence of the randomness can be expected to decreases with $n$, and the expectation of the Brier score can be seen as the long term average, which represent typical values if $n$ is big. More details on properties as well as calculations for the derivations above can be found in Appendix §$\S$ A, and connections to other measures in Appendix §$\S$ C.
2.1. Means of understanding the Brier score
  • Expectation of the Brier score: We will analyze the expected value the Brier score (1.1) takes, which will help to illustrate typical observed values of the Brier score.
  • Simulation-Based Evaluations: We assess observations of the Brier score by simulating Bernoulli outcomes under different sampled probability distributions for $q_{i}$ and different sample sizes. More detail is in Appendix §$\S$ F.
    • Sample size $n$:. Settings considered: $n\in\{300,1000\}$.
    • True distribution: We consider $q_{i}$ as realisations of random variables. The $q_{i}\in[0,1]$ are then used to simulate $Y_{i}\sim Bern(q_{i})$.
    • Predictor distribution: The $p_{i}\in[0,1]$ used in (1.1) are considered as functions of $q_{i}$ with potentially random error.
    • Estimand: Includes median Brier score, $5\%,95\%$ quantiles, and the distribution via violin plots.
3. Misconceptions
  • A Brier score of 0 means a perfect model. A Brier score of 0 implies perfect alignment between predicted probabilities and observed outcomes, with predictions exclusively 0 or 1. The true probabilities are within $[0,1]$, usually expected in $(0,1)$. Hence, rather than signalling a perfect model, an observed Brier score of 0 potentially indicates errors.
  • A perfect model has Brier score 0. With at least one of the true probabilities $q_{i}$ in $(0,1)$, observing a Brier score of 0 with perfect predictions is impossible (see Appendix E). Hence in normal situations perfect models will have a Brier score bigger 0.
To illustrate this and for later reference, we present simulations with perfect predictions, showing that ideal models do not yield a Brier score of 0. In fact, expected Brier scores for a perfect model can be notably high.
Figure 1. Violin plot for simulations with $n=1000$. The left side has perfect predictions, the right $Unif(-0.1,0.1)$ noise added.
Brier scores enable model comparison across datasets, but comparisons can be misleading. We examine three scenarios of increasing risk to illustrate this point.
3.2.1. Comparing Models on the same dataset via Brier score:
When evaluating two models on the same dataset, lower Brier score indicate better fit. There are two potential issues that can prevent the better model to have lower Brier score:
  1. (I)
    It is possible that the worse model has a better score by chance. However, with higher sample sizes this becomes more unlikely.
  2. (II)
    In expectation the Brier score ranks according to the $l_{2}$-distance between prediction $p\in[0,1]^{n}$ to the true probability $q\in[0,1]^{n}$. Changing perfect predictions from $q_{i}$ to $q_{i}+\varepsilon$ or $q_{i}-\varepsilon$ gives the same result in expectation. However, for humans the direction can matter particularly for $q_{i}$ close to zero or one [27]. Consider the example with one observation from [27, §$\S$ 3] where the probability of an event is $P(Y_{1}=1)=\frac{1}{10}$ (i.e., 10% chance), and compare two models:
    • Model 1: $p_{1}=0$ (predicting the event will never occur)
    Model 1 has an expected Brier score of $0.1$, as compared to $0.1125$ for model 2. However, model 1 predicts a zero probability for the observation that has actually a 10% probability, hence 25% might seem better from a humans perspective.
3.2.2. Comparing Brier scores across different datasets with similar incidences (also known as class imbalance) is meaningful.
Some recommend to compare Brier score of prediction models on datasets with similar class imbalances [10]. However, the true probability distribution is unobservable, and true values $q_{i}$ strongly influences the expectation of the Brier score. Thus, even with identical class imbalances, we cannot assume that the true probability distributions are comparable and the same caution as in the previous point should be exercised in interpreting such comparison. As an example, compare three cases: True distribution always $1/2$, independent $Unif(0,1)$ from Figure 1, or a 00 or $1$ distribution with each 50 %. The perfect predictions have in each case expected Brier score $0.25,0.16$ and 00. Hence perfect predictions for true distribution always $1/2$ and expected Brier score $0.25$ is better than, say predictions with noise $Unif(-0.1,0.1)$ for true probabilities $Unif(0,1)$ from Figure 1 with expected Brier score $0.17$. Potentially, similarities in outcome and population trough population characteristics might indicate how true probabilities differ or align in a clinical prediction model setting.
3.2.3. Comparing Brier scores across datasets with different incidences:
If true outcome distributions differ, even perfect models yield different Brier score distributions and (potentially) expectations. Thus, cross-dataset comparisons may lack meaningful insight. Observing different incidences may indicate that the true outcome distributions differ, making the Brier score comparison unreliable. Compare, e.g. the $Beta(1,3)$ prediction with prediction error $+Unif(-0.1,0.1)$ to the perfect smoke prediction model from figure 1. The $Beta(1,3)$ prediction with prediction error has lower Brier score mostly when compared to perfect smoking prediction, nonetheless the perfect model is obviously better.
3.3. Misconception #3: A low Brier score indicates good calibration.
A low Brier score does not necessarily indicate good calibration of a model. We can have perfect predictions, but low or high Brier score due to the underlying probabilities, or similary not so accurate or biased predictions but a Brier score that is low or high due to the underlying probabilities. In expectation a change of perfect prediction from $p_{i}=q_{i}$ to $q_{i}+\varepsilon$ or $q_{i}-\varepsilon$ gives the same result (cf., e.g., (A.4)), and miscalibration where errors tend to go mostly in one direction are equally punished, but e.g. calibration is differently affected as illustrated in Figure 2.
Assessing calibration should be done using additional metrics like calibration in the large (CIL), calibration curves or similar evaluation components [6, 39]. Hence Brier score should not be the sole criterion for evaluating model performance.
Figure 2. Violin plot for simulations with $n=300$.
Note that the lowest expected Brier score occurs when predicted probabilities match the true probabilities, representing a perfect prediction that cannot be improved. However, we cannot observe the true probabilities. With observed incidences of $0.5$, perfect predictions can yield a Brier score of close to 0.25. Hence if we observe a Brier score of around $\bar{y}-\bar{y}^{2}$, the following alternatives to a bad model could explain such a Brier score:
  • Many of the true probabilities are around $\bar{y}$, making expected Brier scores of perfect predictions close to $\bar{y}-\bar{y}^{2}$.
  • For $n$ low, randomness can make the Brier score higher than its expectation.
As an illustration of example values, consider $\bar{y}-\bar{y}^{2}-BS_{perf}$ in figure 3 on the left, where $BS_{perf}$ is the Brier score under perfect predictions. The observed median is very low, with $5\%$ percentile below or around zero.
Figure 3. Violin plot for simulations with $n=300$.
Although the observed Brier score might often be bounded above by $\bar{y}-\bar{y}^{2}$ in practice, higher values can occur. Since true probabilities $q_{i}$ are unobservable, we cannot rule out that even perfect predictions yield a score near the unobservable $\bar{q}-\bar{q}^{2}$, where $\bar{q}$ is the true mean incidence $\bar{q}=\frac{1}{n}\sum_{i=1}^{n}q_{i}$. The Brier score’s randomness stems from the outcomes, also making $\bar{y}$ random. Hence a Brier score exceeding $\bar{y}-\bar{y}^{2}$ may also result from the same options as given in misconception #4. To illustrate this, we show for perfect predictions in some settings with $n=300$, the probability of having a Brier score that is bigger than $\bar{y}-\bar{y}^{2}$ is nonzero. As in practice we will not have perfect predictions, for reasonable predictions the probability to have a Brier score bigger than $\bar{y}-\bar{y}^{2}$ is higher.,The right part can be understood as lower bounds for corresponding probabilities in the corresponding settings.
Figure 4. Histogram for simulations with $n=300$.
4. Conclusions and final remarks
We adressed common misconceptions regarding the interpretation and use of the Brier score. The Brier score is a realisation of a random variable. The true underlying probabilities influence the expectation of Brier score strongly, potentially stronger than the closeness of the predictions to the true probabilities. Brier scores of zero are rather an indication of errors in realistic settings, and a low Brier score does not necessarily indicate a perfect model. Comparisons of Brier scores across models on the same data can be done, and across different data should be avoided or interpreted carefully. Low Brier scores do not guarantee good calibration as evaluating calibration should be done with other metrics. In analogy to idealised settings, randomness in observed Brier scores can be expected to decrease with bigger sample sizes. A recent literature review on clinical prediction models found that sample sizes used had median sample size of $1250$ with $(Q1,Q3)=(353,188860)$[13]. Hence at least a quarter of the prediction models had relatively low sample sizes, with randomness similar to the settings in our simulations with $n=300$ or $n=1000$.
The Brier score remains a valuable metric for assessing probabilistic predictions, but its interpretation requires an understanding of how the underlying probabilities and closeness of predictions influence the observed value. Once misconceptions are avoided, the Brier score serves as a reliable relative measure of overall performance. In particular, it is effective and strictly proper and hence in expectation, it reflects Euclidean distance between predictions, and is minimised uniquely at the true probabilities making relative comparisons on the same data meaningful.
Appendix A More on the Brier score
@@
In this section, we review key properties of the Brier score, delve into some mathematical points, and give more details on the simulations we use to visualise the behavior of the Brier score.
When evaluating the Brier score (1.1), one can compare (1.1) to the trivial Brier score when entering $p_{1/2}=(1/2,\cdots,1/2)$, which, independent of the specific value of $y$ gives $BS(p_{1/2},y)=1/4$. A slightly better prediction is the mean incidence $\bar{y}_{v}=(\bar{y},\cdots,\bar{y})$, which when used in the Brier score gives
@@
Clearly, $BS(p_{1/2},y)=1/4\geq BS(\bar{y}_{v},y)=\bar{y}-\bar{y}^{2}$. For high or low incidences, $BS(\bar{y}_{v},y)$ is low, and, e.g., for $\bar{y}=0.1$ or $\bar{y}=0.9$, $BS(\bar{y}_{v},y)=0.09$. Furthermore, $\bar{y}-\bar{y}^{2}$ is symmetric around $0.5$, see Figure 5.
A.1. Key properties about the Brier score
We summarise essential properties of the Brier score.